OpenStudy (chillhill):

Can anyone please explain this to me: (x+2/x-1)=x/2 I know you cross multiple by it says the answer is -1 or 4 and I'm not sure how to get these answers

jimthompson5910 (jim_thompson5910):

you should have this when you cross multiply 2(x+2) = x(x-1) 2x + 4 = x^2 - x 0 = x^2 - x - 2x - 4 x^2 - 3x - 4 = 0